Calculate the heat energy released when 22.3 g of liquid mercury at 25.00 C is converted to solid mercury at its melting point. Number Constants for mercury at 1 atm heat capacity of Hg() 28.0 J/(mol. K) 234.32 K melting point enthalpy of fusion 2.29 kJ/mol
Answer #1

Molar mass of mercury = 200.592 give/mol. No.of moles is equal to 22.3/200.592 = 0.11117 mol,. H = 0.11117. * 28 *(298-234.32) = 198.22 J,. 234.32K liquid to 234.32 K solid then H. =0.11117. * 2.29. = 0.2544 KJ total heat = 0.1982. +0.2544 = 0.4532 KJ.